OpenStudy (anonymous):

x^2=2(1-k)x-9-k^2 has 2 real roots.Find the range of k. pls check my working and answer whether is correct or not. @ganeshie8

OpenStudy (anonymous):

\[x^2=2x-2kx-9-k^2\]\[x^2-2x+2kx+9+k^2=0\]\[x^2-x(2-2k)+9+k^2=0\]\[b^2-4ac>0\]\[(2-2k)^2-4(1)(9+k^2)>0\]\[(2-2k)(2-2k)-36-4k^2>0\]\[4-8k+4k^2-36-4k^2>0\]\[-8k-32>0\]\[-8k>32\]\[k<-4\]

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

Looks perfect!

OpenStudy (anonymous):

Thnx @ganeshie8